# A Statistical Approach for Synthetic EEG Data Generation

### v1

@@ 4 Results
Figure 10: Distribution analysis of original vs. synthetic data generated using a VAE.
5 Discussion
The primary objective of this study was to develop and validate a statistical approach capable of producing high-quality synthetic EEG data. To achieve this, we utilized four original datasets that incorporated additional biomarkers, emotion scores, and stress levels recorded during complex task performance. These diverse datasets provided a robust foundation for experimentation and model development. Our synthesis approach employed a hybrid methodology [17] combining random sampling from the original datasets with Spearman correlation-based filtering to ensure the desired EEG duration and feature consistency.
Validation of the generated synthetic EEG data was conducted using a comprehensive 5-step quality evaluation framework. This framework encompassed distribution analysis, statistical testing, and the application of machine learning classifiers to rigorously assess the fidelity of the synthetic data in representing the original datasets. The results demonstrated that the synthetic EEG data preserved the key characteristics and variability of the original data along with non-EEG biomarkers, underlying emotion and acute stress response, making it suitable for downstream applications such as research, algorithm development, and training of machine learning models.
A notable advantage of our method is its simplicity and efficiency, requiring significantly lower computational resources compared to more complex approaches such as GANs or diffusion models. This efficiency is particularly advantageous in scenarios where computational resources are limited or rapid data generation is required. Furthermore, the cost-effectiveness of synthetic EEG data is underscored by the high expense of acquiring real EEG recordings, which can exceed \$800 per session [11]. By utilizing synthetic data, researchers can substantially reduce research costs while maintaining access to diverse and representative datasets.
In addition to cost savings, the use of synthetic EEG data offers significant benefits in terms of patient privacy. Real EEG data is inherently sensitive, and its use in research and machine learning carries potential risks related to data security and ethical concerns. Synthetic data mitigates these risks by eliminating direct links to individual patients, thereby enhancing privacy while enabling wide-ranging applications. The proposed method further highlights the potential of hybrid generative approaches to address the growing demand for accessible, high-quality EEG data. By advancing synthetic data methodologies, this research contributes to the broader goal of enabling ethical, cost-effective, and scalable solutions for neuroscience and related disciplines.
@@ 7 Conclusion
This study introduces a scalable and cost-efficient method for generating synthetic EEG data, addressing critical challenges such as the high costs of traditional EEG acquisition and the limited availability of open-access datasets. Synthetic EEG data offers a transformative opportunity to overcome these barriers, enabling the development of robust and privacy-preserving datasets that facilitate the training of machine learning models for healthcare applications. By employing correlation analysis and random sampling, the proposed approach produces synthetic datasets that closely replicate the statistical and structural properties of real-world EEG data. The inability of machine learning models to differentiate synthetic samples from the original, underscores the high fidelity of the generated data. To encourage further exploration and application of this method, the complete source code used in this study is publicly available on GitHub at https://github.com/xalentis/SyntheticEEG.

### v2

@@ 4 Results
Figure 10: Distribution analysis of original vs. synthetic data generated using a VAE. Substantial variation occurs across all frequencies between the original and synthetically generated samples.
@@ 6 Conclusion
This study presents a novel, scalable, and cost-efficient method for generating synthetic EEG data that addresses key limitations of traditional EEG acquisition including acquisition cost and public data availability. By combining correlation analysis with random sampling, the approach produces synthetic EEG data that closely replicate the statistical and structural properties of real signals, including complex neuro-physiological relationships tied to emotion and stress.
Validation was performed using diverse datasets incorporating EEG alongside complementary biomarkers and emotion scores, with a rigorous five-step evaluation framework encompassing statistical tests and classifier-based assessments. This comprehensive validation process confirmed that the synthetic data preserved essential signal characteristics, variability, and meaningful physiological patterns relevant to brain function under stress and cognitive load.
Beyond technical advantages, synthetic EEG data generated via this approach substantially enhances patient privacy by eliminating direct links to individual subjects, thus mitigating ethical and security concerns associated with real EEG data. By delivering an accessible, validated, and privacy-preserving data generation solution, this work advances the field toward scalable and ethical EEG data augmentation, ultimately empowering broader research and machine learning applications across neurological and mental health domains. To encourage further exploration and application of this method, the complete source code used in this study is publicly available on GitHub at https://github.com/xalentis/SyntheticEEG.